Please help me understand this one. Evaluate exponential function for given x value.
f(x) = 2^x , x=3.2 . Thanks.

Answer by Nate(3500)

f(x) = 2^x
f(x) = 2^(3.2)
f(x) = 2^(16/5)
f(x) = 2*2*2*(2^(1/5))
f(x) = 8*(2^(1/5)) that means fifth root of two
f(x) = 9.1895868....